OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192.

OpenStudy (anonymous):

a(n)=a(1) * r^ (n-1) thus, a(6) = a(1) * r^5 -8192= 8*r^5 r=4.

OpenStudy (anonymous):

wat formula do i plug r into to get my answer

OpenStudy (anonymous):

The first one, that's the general formula for a GP.

OpenStudy (anonymous):

i keep getting 682.66...

OpenStudy (anonymous):

Thats too low!

OpenStudy (anonymous):

A) 134,217,728 B) 33,554,432 C) -33,554,432 D) -134,217,728 those are the answer choices

OpenStudy (anonymous):

Yup, that has the answer.

OpenStudy (anonymous):

so is the answer b...

OpenStudy (anonymous):

Nope, Sorry I did a mistake you need to figure out in my first post. A simple calculation mistake!

OpenStudy (anonymous):

thanks ill try to figure it out from there then :)

OpenStudy (anonymous):

Yeah! Its my pleasure! :)